# Foundation Inference Models for Markov Jump Processes

### v2

@@ 3.2 Supervised Inference Models
We first process each time series with a network $\psi_{1}$ to get a set of $K$ embeddings, which we then summarize into a global representation $\hat{h}_{\theta}$ through the attention network $\Omega_{1}$. In equations, we write
$\hat{h}_{\theta}=\Omega_{1}(h_{1\theta},\dots,h_{K\theta},\theta)\,\,\,\text{with}\,\,\,h_{k\theta}=\psi_{1}(x_{k1}^{\prime},\tau_{k1},\dots,x_{kl}^{\prime},\tau_{kl},\theta).$ (4)
Next we use the global representation to get an estimate of the intensity rate matrix, which we model as a Gaussian variable, and the initial distribution of the hidden MJP as follows
$\mathbf{\hat{F}}=\phi_{1}(\hat{h}_{\theta},\theta),\quad\log\text{Var}\,\mathbf{\hat{F}}=\phi_{2}(\hat{h}_{\theta},\theta),\,\,\,\text{and}\quad\hat{\pi}_{0}=\phi_{3}(\hat{h}_{\theta},\theta).$ (5)
Figure 2 summarizes the inference model.
@@
$\mathcal{L}$ $=$ $-\underset{\mathbf{F},\mathbf{A}\sim p_{{\text{rates}}}}{\mathbb{E}}\Big{\{}\sum_{ij}a_{ij}\Big{[}\frac{(f_{ij}-\hat{f}_{ij})^{2}}{2\text{Var}\hat{f}_{ij}}+\frac{1}{2}\log\text{Var}\hat{f}_{ij}\Big{]}-\lambda(1-a_{ij})\Big{[}\hat{f}_{ij}^{2}+\text{Var}\,\hat{f}_{ij}\Big{]}\Big{\}}$ (6)
$\quad\quad\quad-\underset{\pi_{0}\sim p}{\mathbb{E}}\Big{\{}\sum_{i}\pi_{i0}\log\hat{\pi}_{i0}\Big{\}},$
where the second term is nothing but the mean-squared error of the predicted rates $\hat{f}_{ij}$ (and its standard deviation) when the corresponding link is missing, and can be understood as a regularizer with weight $\lambda$.
FIM context number. During training, FIM processes a variable number $K$ of time series, which lies on the interval $[K_{\text{\tiny min}},K_{\text{\tiny max}}]$. Similarly, each one of these time series has a variable number $l$ of observation points, which lies on the interval $[l_{\text{\tiny min}},l_{\text{\tiny max}}]$. We shall say that FIM needs a bare minimum of $K_{\text{\tiny min}}l_{\text{\tiny min}}$ input data points to function. Perhaps unsurprisingly, we have empirically seen that FIM perform bests when processing $K_{\text{\tiny max}}l_{\text{\tiny max}}$ data points. Going significantly beyond this number seems nevertheless to decrease the performance of FIM. We invite the reader to check Appendix C for details.
Let us define then, for the sake of convenience, the FIM context number $c(K,l)=Kl$ as the number of input points [^2] We can think about it as the context length in large language models. FIM makes use of to estimate $\mathbf{F}$ and $\pi_{0}$.
@@
4 Experiments
In this section we test our methodology on five datasets of varying complexity, and corrupted by noise signals of very different nature, whose hidden MJPs are known to take values in state spaces of different sizes. In what follows we use one and the same (pretrained) FIM to infer hidden MJPs from all these datasets. Our FIM was trained on a dataset of 45K MJPs, defined over state spaces whose sizes range from 2 to 6. A maximum of ($K=$)300 realizations (paths) per MJP were observed during training, everyone of which spanned a time-horizon $T=10$, recorded at a maximum of 100 time points, 1% of which were mislabeled. Given these specifications, FIM is expected to perform best for the context number $c(300,100)$ during evaluation. Additional information regarding model architecture, hyperparameter selection and other training details can be found in Appendix C.
@@ 4.1 The Discrete Flashing Ratchet (DFR): A Proof of Concept
In statistical physics, the ratchet effect refers to the rectification of thermal fluctuations into directed motion to produce work, and goes all the way back to Feynman (Feynman et al., 1965). Here we consider a simple example thereof, in which a Brownian particle, immersed in a thermal bath at unit temperature, moves on a one-dimensional lattice. The particle is subject to a linear, periodic and asymmetric potential of maximum height $2V$ that is switched on and off at a constant rate $r$. The potential has three possible values when it is switched on, which correspond to three of the states of the system, and the particle jumps among them with rate $f_{ij}^{\text{\tiny on}}$. When it is switched off, the particle jumps freely with rate $f_{ij}^{\text{\tiny off}}$. We can therefore think of the system as a six-state system, as illustrated in Figure 4. Similar to Roldán and Parrondo (2010), we now define the transition probability rates as
@@
The results show that FIM performs on par with (or even better than) NeuralMJP, despite not having been trained on the data. Note in particular that our results are sharply peaked around their mean, indicating that a context of $c(300,50)$ points only contains enough information to describe the data well. What is more, Table 13 in the Appendix demonstrates that FIM can infer vanishing transition rates as well. Now, being able to infer the rate matrix in zero-shot mode allows us to immediately estimate a number of observables of interest without any training. Stationary distributions, relaxation times and mean first-passage times (see e.g. Appendix A), as well as time-dependent moments, can all be computed zero-shot via FIM. On the left block of Figure 5 we report, for example, the time-dependent class probabilities (i.e. the master eq. solutions) as computed from the FIM-inferred rate matrix (black) against the ground-truth solution (blue). The agreement is very good.
@@
Zero-shot estimation of entropy production. The DFR model is interesting because the random switching combined with the asymmetry in the potential make it more likely for the particle to jump towards the right (see Figure 5). Indeed, that is the ratchet effect. As a consequence, the system features a stationary distribution with a net current — the so-called non-equilibrium steady state (Ajdari and Prost, 1992), which is characterized by a non-vanishing (stochastic) entropy production. The development of (neural) estimators of entropy production is an active topic of research (see e.g. Kim et al. (2020) and Otsubo et al. (2022)). Given that the entropy production can be written down in closed form as a function of both the rate matrix and the master eq. solution (see e.g. Seifert (2012)), we can readily use our FIM to estimate it.
@@
Figure 5 displays the total entropy production computed with FIM for a set of different potentials. The results are averaged over 15 FIM evaluations with $c(300,50)$ and are again in very good agreement with the ground truth. It is noteworthy that FIM, trained on our heuristically constructed dataset, captures well a continuous set of MJPs. That is, we evaluate one and the same FIM over different datasets, each sampled from a DFR model with a different potential value. In sharp contrast, state-of-the-art models need to be retrained for every new potential value.
@@
Zero-shot simulation of the DFR process. Inferring the rate matrix and initial condition of a MJP process entails that one can also sample from it. Our FIM could thus, in principle, be used as a zero-shot generative model for MJPs. However, to test the quality of said MJP realizations, wrt. some target MJP, we would need a distance between the two. Here we propose to use the Hellinger distance (Le Cam and Yang, 2000) to first estimate the divergence between a sequence of (local) histogram pairs, recorded at a given set of observation times, and then average the local estimates along time. Appendix E.1 empirically demonstrates that this pragmatically defined MJP distance is sensible.
@@
Table 7 reports the time-averaged Hellinger distance between 1000 (ground-truth) DFR paths and 1000 paths sampled from (the MJPs inferred by) NeuralMJP and FIM. We repeat this calculation 100 times, for 1000 newly sampled paths from NeuralMJP and FIM, but the same 1000 target paths, to compute the mean values and error bars in the Table. The results show that the zero-shot DFR simulation obtained through FIM is on par to the NeuralMJP-based simulation wrt. the ground truth.
@@ 4.2 Switching Ion Channel (IonCh): Zero-Shot Inference of Three-State MJP
In this section we study the conformational dynamics of the viral ion channel $\text{Kcv}_{\tiny\text{MT325}}$ which exhibits three metastable states (Gazzarrini et al., 2006). Specifically, we analyse the ion flow across the membrane as the system jumps between its metastable configurations, which was recorded at a frequency of 5kHz over one second. Figure 1 shows one snapshot of these recordings, which were made available to us via private communication. Our goal is to infer physical observables, like the stationary distribution and the mean first-passage time of the conformational dynamics, and to compare our findings against the SDiff model of Köhs et al. (2021) and NeuralMJP.
The measurements live in real space, which implies that we need to obtain a coarse-grained representation (CGR) out of them. Here we consider two: the CGR inferred by NeuralMJP and a naive CGR obtained with a Gaussian Mixture Model (GMM). Since we only have 5000 observations available, we make use of a single FIM evaluation with context number $c(50,100)$. We infer two FIM rate matrices, one per each CGR. Let us label them as FIM-NMJP and FIM-GMM.
Table 7 contains the inferred stationary distributions from all models and evidences that a single FIM evaluation is enough to unveil the long-time asymptotics of the process. Similarly, Table 12 in the Appendix, which contains the inferred mean-first passage times, demonstrates that FIM makes the same inference about the short-term dynamics of the process as do SDiff and NeuralMJP without any training. See Appendix E for additional results.
@@ 4.3 Alanine Dipeptide (ADP): Zero-Shot Inference of Six-State MJP
Alanine dipeptide is 22-atom molecule widely used as a benchmark in molecular dynamics simulation studies. The main reason being that the heavy-atom dynamics, which jumps between six metastable states, can be fully described in terms of the dihedral (torsional) angles $\psi$ and $\phi$ (Mironov et al., 2019).
We examine an all-atom ADP simulation of 1 microsecond and compare against both, the VampNets model of Mardt et al. (2017) and NeuralMJP. The data consists of the values taken by the dihedral ADP angles as time evolves and thus needs to be mapped onto some coarse-grained space. We again make us of NeuralMJP to obtain a CGR. We then use FIM with context number $c(300,100)$ to process 32 100-point time windows of the simulation and compute an average rate matrix. Note that this is the optimal context number of our pretrained model. Table 1 (and Appendix E.2) confirms that, once again, FIM can infer the same physical properties from the ADP simulation as the baselines.
Zero-shot simulation of the alanine dipeptide. Simulations in coarse-grained space for molecular dynamics is a high-interest research direction (Husic et al., 2020). Here we demonstrate that FIM can be used to simulate the ADP process in zero-shot mode. Indeed, Table 7 reports the distance from both NeuralMJP and FIM to a target ADP process, computed from 200 paths with 100 observations each. Once more, FIM performs comparable to NeuralMJP
@@ 4.4 Zero-Shot Inference of Two-State MJPs
Finally, we consider two additional system that feature jumps between two metastable states: a simple protein folding model and a two-mode switching system. We invite the reader to check out Appendix E.5 and E.6 for details. That being said, Table 11 reports the distance of both NeuralMJP and FIM wrt. the empirical protein folding process (PFold). The high variance indicates that the distance cannot resolve any difference between the processes given the available number of samples.
@@ 5 Conclusions
In this work we introduced a novel methodology for zero-shot inference of Markov jump processes and its Foundation Inference Model (FIM). We empirically demonstrated that one and the same FIM can be used to estimate stationary distributions, relaxation times, mean first-passage times, time-dependent moments and thermodynamic quantities (i.e. the entropy production) from noisy and discretely observed MJPs, taking values in state spaces of different dimensionalities, all in zero-shot mode. FIM is also the first zero-shot generative model for MJPs. Future work involves extending our methodology to Birth and Death processes, and Point processes in particular.
@@ 6 Limitations
The main limitations of our methodology clearly involves our synthetic training dataset. Evaluating FIM on empirical datasets whose distribution deviates significantly from our synthetic distribution will inevitably yield poor estimates. Consider Figure 5 (right), for example. The performance of FIM quickly deteriorates for $V\geq 3$, for which the ratio between the largest and smallest rates gets larger than about three orders of magnitude. These cases lie out of our synthetic distribution.

### v3

@@ 3.2 Supervised Recognition Model
We first process each time series with a network $\psi_{1}$ to get a set of $K$ embeddings, which we then summarize into a global representation $\mathbf{h}_{\theta}$ through the attention network $\Omega_{1}$. In equations, we write
$\mathbf{h}_{\theta}=\Omega_{1}(\mathbf{h}_{1\theta},\dots,\mathbf{h}_{K\theta},\theta)\,\,\,\text{with}\,\,\,\mathbf{h}_{k\theta}=\psi_{1}(x_{k1}^{\prime},\tau_{k1},\dots,x_{kl}^{\prime},\tau_{kl},\theta)\,\,\text{and}\,\,k=1,\dots,K.$ (4)
Next we use the global representation to get an estimate of the intensity rate matrix, which we artificially model as a Gaussian variable with positive mean, and the initial distribution of the hidden MJP as follows
$\mathbf{\hat{F}}=\exp(\phi_{1}(\mathbf{h}_{\theta},\theta)),\quad\text{Var}\,\mathbf{\hat{F}}=\exp(\phi_{2}(\mathbf{h}_{\theta},\theta))\,\,\,\text{and}\quad\boldsymbol{\hat{\pi}}_{0}=\phi_{3}(\mathbf{h}_{\theta},\theta),$ (5)
where the exponential function ensures the positivity of our estimates, and the variance is used to represent the model’s uncertainty in the estimation of the rates (Seifner et al., 2024). The right panel of Figure 2 summarizes the recognition model, and Appendix C provides additional information about the inputs to, outputs of and rescalings done by the model.
@@
$\mathcal{L}$ $=$ $-\underset{\mathbf{F},\mathbf{A}\sim p_{{\text{rates}}}}{\mathbb{E}}\Big{\{}\sum_{ij=1}^{C}a_{ij}\Big{[}\frac{(f_{ij}-\hat{f}_{ij})^{2}}{2\text{Var}\hat{f}_{ij}}+\frac{1}{2}\log\text{Var}\hat{f}_{ij}\Big{]}-\lambda(1-a_{ij})\Big{[}\hat{f}_{ij}^{2}+\text{Var}\,\hat{f}_{ij}\Big{]}\Big{\}}$ (6)
$\quad\quad\quad-\underset{\boldsymbol{\pi}_{0}\sim p}{\mathbb{E}}\Big{\{}\sum_{i=1}^{C}\pi_{i0}\log\hat{\pi}_{i0}\Big{\}},$
where the second term is nothing but the mean-squared error of the predicted rates $\hat{f}_{ij}$ (and its standard deviation) when the corresponding link is missing, and can be understood as a regularizer with weight $\lambda$. The latter is a hyperparameter.
FIM context number. During training, FIM processes a variable number $K$ of time series, which lies on the interval $[K_{\text{\tiny min}},K_{\text{\tiny max}}]$. Similarly, each one of these time series has a variable number $l$ of observation points, which lies on the interval $[l_{\text{\tiny min}},l_{\text{\tiny max}}]$. We shall say that FIM needs a bare minimum of $K_{\text{\tiny min}}l_{\text{\tiny min}}$ input data points to function. Perhaps unsurprisingly, we have empirically seen that FIM perform bests when processing $K_{\text{\tiny max}}l_{\text{\tiny max}}$ data points. Going significantly beyond this number seems nevertheless to decrease the performance of FIM. We invite the reader to check Appendix D for details.
Let us define then, for the sake of convenience, the FIM context number $c(K,l)=Kl$ as the number of input points [^4] We can think about it as the context length in large language models. FIM makes use of to estimate $\mathbf{F}$ and $\boldsymbol{\pi}_{0}$.
4 Experiments
In this section we test our methodology on five datasets of varying complexity, and corrupted by noise signals of very different nature, whose hidden MJPs are known to take values in state spaces of different sizes. In what follows we use one and the same (pretrained) FIM to infer hidden MJPs from all these datasets, without any parameter fine-tuning. Our FIM was (pre)trained on a dataset of 45K MJPs, defined over state spaces whose sizes range from 2 to 6. A maximum of ($K=$)300 realizations (paths) per MJP were observed during training, everyone of which spanned a time-horizon $T=10$, recorded at a maximum of 100 time points, 1% of which were mislabeled. Given these specifications, FIM is expected to perform best for the context number $c(300,100)$ during evaluation. Additional information regarding model architecture, hyperparameter selection and other training details can be found in Appendix D.
All these baselines are trained on the target datasets.
@@ 4.1 The Discrete Flashing Ratchet (DFR): A Proof of Concept
In statistical physics, the ratchet effect refers to the rectification of thermal fluctuations into directed motion to produce work, and goes all the way back to Feynman (Feynman et al., 1965). Here we consider a simple example thereof, in which a Brownian particle, immersed in a thermal bath at unit temperature, moves on a one-dimensional lattice. The particle is subject to a linear, periodic and asymmetric potential of maximum height $2V$ that is switched on and off at a constant rate $r$. The potential has three possible values when is switched on, which correspond to three of the states of the system. The particle jumps among them with rate $f_{ij}^{\text{\tiny on}}$. When the potential is switched off, the particle jumps freely with rate $f_{ij}^{\text{\tiny off}}$. We can therefore think of the system as a six-state system, as illustrated in Figure 4. Similar to Roldán and Parrondo (2010), we now define the transition rates as
@@
The results show that FIM performs on par with (or even better than) NeuralMJP, despite not having been trained on the data. Note in particular that our results are sharply peaked around their mean, indicating that a context of $c(300,50)$ points only contains enough information to describe the data well. What is more, Table 13 in the Appendix demonstrates that FIM can infer vanishing transition rates as well (see Eq. 6). Now, being able to infer the rate matrix in zero-shot mode allows us to immediately estimate a number of observables of interest without any training. Stationary distributions, relaxation times and mean first-passage times (see Appendix A for their definition), as well as time-dependent moments, can all be computed zero-shot via FIM. For example, we report on the left block of Figure 5 the time-dependent class probabilities (i.e. the master eq. solutions) computed with the FIM-inferred rate matrix (black), against the ground-truth solution (blue). The agreement is very good.
@@
Zero-shot estimation of entropy production. The DFR model is interesting because the random switching combined with the asymmetry in the potential make it more likely for the particle to jump towards the right (see Figure 5). Indeed, that is the ratchet effect. As a consequence, the system features a stationary distribution with a net current — the so-called non-equilibrium steady state (Ajdari and Prost, 1992), which is characterized by a non-vanishing (stochastic) entropy production. The development of (neural) estimators of entropy production is a very active topic of current research (see e.g. Kim et al. (2020) and Otsubo et al. (2022)). Given that the entropy production can be written down in closed form as a function of both the rate matrix and the master eq. solution (see e.g. Seifert (2012)), we can readily use FIM to estimate it.
@@
Figure 5 displays the total entropy production computed with FIM for a set of different potentials. The results are averaged over 15 FIM evaluations with $c(300,50)$ and are again in very good agreement with the ground truth. It is noteworthy that FIM, trained on our heuristically constructed dataset, captures well a continuous set of MJPs. That is, we evaluate one and the same FIM over different datasets, each sampled from a DFR model with a different potential value. In sharp contrast, state-of-the-art models need to be retrained for every new potential value (Kim et al., 2020).
@@
Zero-shot simulation of the DFR process. Inferring the rate matrix and initial condition of a MJP process entails that one can also sample from it. Our FIM can thus be used as a zero-shot generative model for MJPs. However, to test the quality of said MJP realizations wrt. some target MJP, we need a distance between the two. Here we propose to use the Hellinger distance (Le Cam and Yang, 2000) to first estimate the divergence between a sequence of (local) histogram pairs, recorded at a given set of observation times, and then average the local estimates along time. Appendix F.1 empirically demonstrates that this pragmatically defined MJP distance is sensible.
@@
Table 7 reports the time-averaged Hellinger distance between 1000 (ground-truth) DFR paths and 1000 paths sampled from (the MJPs inferred by) NeuralMJP and FIM. We repeat this calculation 100 times, for 1000 newly sampled paths from NeuralMJP and FIM, but the same 1000 target paths, to compute the mean values and error bars in the Table. The results show that the zero-shot DFR simulation obtained through FIM is on par with the NeuralMJP-based simulation, wrt. the ground truth.
In this section we study the conformational dynamics of the viral ion channel $\text{Kcv}_{\tiny\text{MT325}}$, which exhibits three metastable states (Gazzarrini et al., 2006). Specifically, we analyse the ion flow across the membrane as the system jumps between its metastable configurations. This ion flow was recorded at a frequency of 5kHz over one second. Figure 1 shows one snapshot of these recordings, which were made available to us via private communication (see the Acknowledgements). Our goal is to infer physical observables — like the stationary distribution and mean first-passage times — of the conformational dynamics, and to compare our findings against the SDiff model of Köhs et al. (2021) and NeuralMJP.
The recordings live in real space, which means that we first need to obtain a coarse-grained representation (CGR) from them, before we can apply FIM. Here we consider two CGRs: the CGR inferred by NeuralMJP and a naive CGR obtained with a Gaussian Mixture Model (GMM). Given that we only have 5000 observations available, we make use of a single FIM evaluation with context number $c(50,100)$. We infer two FIM rate matrices, one per each CGR, which we label as FIM-NMJP and FIM-GMM.
Table 7 contains the inferred stationary distributions from all models and evidences that a single FIM evaluation is enough to unveil the long-time asymptotics of the process. Similarly, Table 12 in the Appendix, which contains the inferred mean-first passage times, demonstrates that FIM makes the same inference about the short-term dynamics of the process as do SDiff and NeuralMJP. See Appendix F for additional results.
@@ 4.3 Alanine Dipeptide (ADP): Zero-Shot Inference of Six-State MJP
Alanine dipeptide is 22-atom molecule widely used as benchmark in molecular dynamics simulation studies. Its popularity stems from the fact that the heavy-atom dynamics, which jumps between six metastable states, can be fully described in terms of the dihedral (torsional) angles $\psi$ and $\phi$ (see e.g. Mironov et al. (2019) for details).
We examine an all-atom ADP simulation of 1 microsecond, which was made available to us via private communication (see the Acknowledgements below), and compare against both, the VampNets model of Mardt et al. (2017) and NeuralMJP. The data consists of the values taken by the dihedral angles as time evolves and thus needs to be mapped onto some coarse-grained space. We again make use of NeuralMJP to obtain a CGR. We then use FIM with context number $c(300,100)$ to process 32 100-point time windows of the simulation and compute an average rate matrix. Note that this is the optimal context number of our pretrained model. Table 1 (and Appendix F.2) confirms that, once again, FIM can infer the same physical properties from the ADP simulation as the baselines.
Zero-shot simulation of the alanine dipeptide. Simulations in coarse-grained space for molecular dynamics is a high-interest research direction (Husic et al., 2020). Here we demonstrate that FIM can be used to simulate the ADP process in zero-shot mode. Indeed, Table 7 reports the distance from both NeuralMJP and FIM to a target ADP process, computed from 200 paths with 100 observations each. Once more, FIM performs comparable to NeuralMJP.
@@ 4.4 Zero-Shot Inference of Two-State MJPs
Finally, we consider two additional systems that feature jumps between two metastable states: a simple protein folding model and a two-mode switching system. We invite the reader to check out Appendix F.5 and F.6 for the details. That being said, Table 11 reports the distance of both NeuralMJP and FIM wrt. the empirical protein folding process (PFold). The high variance indicates that the distance cannot resolve any difference between the processes given the available number of samples.
@@ 5 Conclusions
In this work we introduced a novel methodology for zero-shot inference of Markov jump processes and its Foundation Inference Model (FIM). We empirically demonstrated that one and the same FIM can be used to estimate stationary distributions, relaxation times, mean first-passage times, time-dependent moments and thermodynamic quantities (i.e. the entropy production) from noisy and discretely observed MJPs, taking values in state spaces of different dimensionalities, all in zero-shot mode. To the best of our knowledge, FIM is also the first zero-shot generative model for MJPs.
Future work shall involve extending our methodology to Birth and Death processes, as well as considering more complex (prior) transition rate distributions. See our discussion on Limitations in the next section, for details.
@@ 6 Limitations
The main limitations of our methodology clearly involve our synthetic distribution. Evaluating FIM on empirical datasets whose distribution significantly deviates from our synthetic distribution will, inevitably, yield poor estimates. Consider Figure 5 (right), for example. The performance of FIM quickly deteriorates for $V\geq 3$, for which the ratio between the largest and smallest rates gets larger than about three orders of magnitude. These cases are unlikely under our prior Beta distributions, and hence effectively lie outside of our synthetic distribution.
More generally, the MJP dynamics underlying phenomena that feature long-lived, metastable states, ultimately depends on the shape of the energy landscape characterizing the set $\mathcal{X}$, inasmuch as the transition rates between metastable states $i$ and $j$ ($f_{ij}$ in our notation) are characterized by the depth of the energy traps (that is, the height of the barrier between them).
In equations, we write
$f_{ij}=\exp\left(\frac{-E_{j}}{T}\right),$ (8)
where $E_{j}$ is the $j$th trap depth, and $T$ is the temperature of the system. Therefore, the distribution over energy traps determines the distribution over transition rates.
Just to give an example, if we studied systems with exponentially distributed energy traps — as e.g. in the classical Trap model of glassy systems of Bouchaud (1992) — we would immediately find $p(f)\propto Tf^{T-1}$. Transition rates sampled from such power-law distributions clearly lie outside our ensemble of Beta distributions, even if we use our rescaling trick. Future work shall explore training FIM on synthetic MJPs featuring power-law-distributed transition rates.